# Radiation forces and torques in optics and acoustics

### v1

@@ V.1.2 Acoustics
These equations are similar to electromagnetic Eqs. (49) with and , i.e., without polarization degrees of freedom. Therefore, they result in similar basic forces (22) and torque (23) or (26) on an isotropic Rayleigh particle: (i) the -directed gradient force, (ii) the -directed radiation-pressure force, and (iii) the -directed torque due to the transverse spin . Such forces and torque were recently described in [543, 14, 203, 461].
The recoil force (24) is produced by the complex kinetic momentum density . In contrast to the optical case (50), it only has and components, and, hence, presents only small corrections to the main gradient and radiation-pressure forces. In acoustics, a recoil force with nontrivial direction requires either anisotropic-shape particles [513] or Willis coupling [482].
Notably, the evanescent-wave example is not only important per se but also as a useful model for all cases where the wave intensity and phase gradients have orthogonal directions. For example, an interference of two plane waves with equal amplitudes and frequencies, but different propagation directions, produces a field with the mean wavevector (phase gradient) directed along, say, the -axis, and with periodic intensity gradients in the orthogonal -direction. Such field produces similar local forces and torques, both in optics [69, 139, 4, 524] and acoustics [488, 103].
@@ V.2 Vortex beams
Let us consider another important example of structured wavefields: paraxial vortex beams. Such beams carrying orbital angular momentum (OAM) were introduced in optics in the beginning of the 1990s [66, 16], and since then they have found numerous applications in optics [15, 547, 68, 21, 386, 210, 601, 486], acoustics [259, 327, 250, 136], and quantum electron/neutron/atomic waves [94, 554, 95, 351, 317, 151, 356]. Independently, of their nature, paraxial vortex beams are characterized by the scalar wave vortex factor , where are cylindrical coordinates, is the radial function which behaves as for , and is the integer topological charge (azimuthal quantum number) of the vortex.
@@ V.2.1 Optics
where is the same complex polarization parameter as we used for evanescent waves, and we neglected the small longitudinal -components of the fields. Substituting the fields (53) into Eqs. (2)–(6) yields the energy, momentum, and spin density distributions:
@@
Here, the azimuthal component of the canonical momentum, , generates the OAM density , which looks similar to the spin but with the vortex charge instead of the helicity number .
@@
Substituting Eqs. (54) into the main force and torque Eqs. (22), (23), and (26), one can see that an isotropic absorbing Rayleigh particle in a paraxial vortex beam experiences: (i) a radial gradient force, (ii) a longitudinal and azimuthal radiation-pressure force, and (iii) a longitudinal torque due to the spin. Therefore, when the -position of the particle is fixed (e.g., by a substrate), it can: (i) be trapped at the radial maximum or minimum of the field intensity by the gradient force (for the positive or negative real part of its dipole polarizability), (ii) orbit azimuthally along a circle due to the azimuthal radiation-pressure force controlled by the vortex charge , and (iii) spin around the local -axis passing through the particle due to the torque controlled by the polarization helicity , see Fig. 16(a,b). The azimuthal orbiting and local spinning demonstrate the qualitative difference between the orbital (extrinsic) and spin (intrinsic) angular momenta. Such behaviour of small particles in optical vortex beams was observed in [215, 216, 409, 223, 558, 224, 161]. Note that the -controlled spinning of the particles in these experiments was due to the particle’s anisotropy and the corresponding anisotropy-induced torque (29) rather than the absorption torque (26). Employing anisotropy rather than absorption is more efficient and allows one to avoid unnecessary heating.
@@
Particles experience ‘negative’ gradient forces and are attracted to low-intensity regions (‘dark-field’ trapping), when the real part of their polarizability is negative. This occurs when the refractive index of the dielectric particle is lower than the index of the surrounding medium [216, 223] or for over-resonant frequencies near polarizability resonances (Fig. 10) [182, 608, 394]. Such negative gradient forces allow trapping negative-polarizability particles in the zero-field centers of vortex beams [475, 507, 410, 242].
@@
When the particle occupies the central area of the vortex beam (e.g., when it is trapped in the center or its size is larger than the typical vortex radius), both the azimuthal radiation force and the torque due to the spin produce similar torques and rotations with respect to the beam/particle center. In this case, the ‘orbital’ and ‘spin’ torques act in similar ways producing spinning of the particle in the vortex center [258, 506], and the OAM-controlled ‘orbital torques’ can be used for local rotational manipulation of particles [598, 32].
@@
Furthermore, beyond the paraxial-field approximation (53) and the simplest gradient and radiation-pressure forces, an additional -dependent orbital rotation of particles can appear in vortex beams. First, this can happen due to the -dependent contribution to the azimuthal canonical momentum (i.e., OAM) in nonparaxial vortex beams [5, 620, 84, 104, 100]. Second, the same effect can appear due to the -dependent azimuthal component of the complex Poynting (kinetic) momentum density and the corresponding recoil force (24) from its real [22, 23, 526] and imaginary [623, 592] parts.
@@ V.2.2 Acoustics
In acoustics, paraxial vortex beams can be described by the scalar pressure field and the corresponding velocity field . This longitudinal sound field produces the energy and canonical momenta densities and entirely similar to electromagnetic Eqs. (54) but with no spin density: . (The transverse spin, present both in optics and acoustics, appears only for nonparaxial fields having both transverse and longitudinal components.) As in optics, the azimuthal component of the canonical momentum is responsible for the OAM carried by the beam.
Thus, paraxial acoustic vortex beams produce only a radial gradient force, as well as a radiation pressure force with longitudinal and azimuthal components. There were a number of experiments demonstrating acoustic ‘orbital torques’ on Mie or larger particles in the center of acoustic vortex beams due to the azimuthal radiation pressure force [559, 511, 170, 24, 50]. The dynamical effects of acoustic vortices have found numerous applications, including trapping, levitation, and holographic tweezers [369, 49, 267, 367, 368, 61], rotation of particles up to 6500 revolutions per second [611], manipulation of cells [63, 617], microbubbles [48] and objects in living bodies [228] (see [250] for a review).
Figure 16(c) shows the results of experimental observations [266] of the local orbital motion of microparticles in acoustic vortex fields. This microparticle velocity field can be associated with the local canonical momentum density or Stokes drift distribution in acoustic vortices [see Section IV.10 and Fig. 14(b)].
@@ V.3.1 Optics
Optical tweezers have been the most prolific implementation of optical forces. They use one or more laser beams, focused into a tight spot to trap and manipulate particles. Optical trapping was proposed and first realised by Ashkin and co-workers [42, 38, 41, 43]. The explanation of the effect depends on the scale of the trapped object, Eq. (17). For particles bigger than several wavelengths, a ray-optics picture is relevant. Namely, large dielectric particles act as lenses, re-directing the momentum of the incident rays, with the corresponding recoil pushing them towards the focus [37, 38]. For particles much smaller than the wavelength, the dipolar regime prevails, and the intensity gradient force in Eq. (22) pushes the particle towards the high-intensity focus. For intermediate, wavelength-order particles, neither model is accurate, and one has to consider higher-order multipoles, Mie scattering calculations, or numerical calculations of forces based on general Eqs. (12).
@@
Importantly, optical trapping can be applied to individual atoms [150, 427, 439], Bose-Einstein condensates [252, 147], biological objects [41, 43], and various nanostructures such as quantum dots, nanotubes, graphene flakes, etc. [362, 417, 529, 282]. Optical tweezers can also be employed for nanofabrication, direct microfluidic control, and assembly [217, 536, 417]. This technique allows one not only to control, but also to measure tiny displacements of the particle, since the light scattering depends on the local intensity, which varies across the focus.
@@
Let us consider the main field configurations used in optical tweezers for small particles. In the simplest case, a single focused beam is used. Then, assuming electric-dipole particles, the trapping gradient force (for ) competes with destabilising radiation-pressure and recoil forces and pushing the particle longitudinally out of the trap, Fig. 17(a). According to Sections III.3 and III.4, the gradient force depends on how tightly the beam is focused and on , while the pressure force is proportional to . For lossless Rayleigh particles with , we have , Eq. (III.3), and the gradient force typically prevails over the pressure one. However, the pressure and recoil forces grow with the particle size and refractive index faster than the gradient force, making it challenging to trap larger dielectric particles or high index particles. This can be overcome using very tight focusing, and/or anti-reflection coatings [108, 281]. In addition, for small particles, the gradient force competes with the Brownian motion, which depends strongly on the surrounding medium and the temperature, and can be an issue for metallic (lossy) nanoparticles that become heated [481]. The polarizability resonances in metallic particles can be advantageously exploited [181, 420].
@@
The next field configuration involves two counter-propagating focused beams with equal amplitudes, polarizations, and other parameters. Such field produces a standing wave with alternating maxima and minima (separated by a half-wavelength distance in the longitudinal direction) of the electric energy density, Fig. 17(b). Importantly, the momentum and energy flux densities vanish in such non-propagating field, and so do the pressure and recoil forces. The counter-propagating beams were first used in the pioneering paper by Ashkin [36, 38] for trapping of large several-wavelength-size particles. The subwavelength standing-wave features are not important for such particles, and the wavelength-averaged intensity gradients provide trapping in the focal zone. For subwavelength Rayleigh particles, the standing-wave structure becomes important, and such dipole particles can be trapped with the gradient force at any of the local field intensity maxima (for ) [609] or minima () [608], as shown in Fig. 17(b). Optical trapping in standing waves becomes particularly important in the context of cavity optomechanics [45].
@@
Optical tweezers may involve complex structured (essentially non-Gaussian) beams. Various trappings in the high-intensity (for ) and low-intensity zones have been exploited in vortex beams, see Section V.2, Fig. 16, and Fig. 17(c). Furthermore, “optical bottle beams”, that generate a dark spot surrounded by brightness in all directions, were also used for ‘dark-field’ trapping [29, 276, 590, 54, 497]. Note that photophoretic forces can play an important role in such ‘dark-field’ trapping with vortex and bottle beams [496, 176, 498, 497, 514]. Diffraction-free Bessel beams [191] were used to trap and manipulate particles over large distances in both single-beam [28] and counter-propagating two-beam [131, 295, 597] regimes. Finally, Airy beams that exhibit ‘transverse acceleration’ upon propagation [79, 510] have also been employed [65, 64].
@@
A breakthrough in optical tweezers was provided by the introduction of multi-functional traps often called holographic tweezers [248]. These are arrays of multiple beams of light with different foci controlled simultaneously and independently, allowing simultaneous multi-point control of many trapped particles [189, 473, 385, 474, 190, 449, 338, 162, 53, 514], see Figs. 1(b), 3(a), and 17(e).
@@
Importantly, optical trapping is easily combined with various torques producing rotations of the trapped particles (either anisotropic or absorptive). This stimulated a great progress in studies of optical torques and their applications [213, 258, 506, 413, 121]. These torque-related effects have a remarkable range of angular velocities: from quasi-static 3D alignments and Hz rotations [420, 545, 586] to MHz and GHz rotations where the tensile strength of the material and quantum effects come into play [27, 309, 272, 8, 450, 241].
@@
Optical trapping and tweezers have found numerous applications in a variety of fields. First, they have become an invaluable tool in biology and biophysics [41, 43, 106, 528, 1, 35, 517, 527, 396, 179, 612, 121, 588], Fig. 17(d). Second, optical trapping has played a crucial role in optical cooling of atoms, molecules, nanoparticles, and larger macroscopic objects [150, 427, 377, 134, 336, 233, 157, 532, 335, 168], leading to applications ranging from quantum interference of macroscopic objects [30, 460] to detection of gravitational waves [31]. Third, optical forces near the centers of optical traps act like harmonic-oscillator potentials, which paves the avenue for various optomechanical applications, both in classical and quantum regimes [45, 378, 135, 241, 280, 235, 533]. Finally, 1D, 2D, and 3D arrays of optically trapped atoms, molecules, or particles can serve for quantum simulators and computers [169, 200, 52, 75, 570, 477, 193, 53], atomic clocks [605], and volumetric displays [514], see Figs. 3(a) and 17(e).
@@ V.3.2 Acoustics
Acoustic trapping and tweezers share natural similarities with their optical counterparts. Nonetheless, there are some important peculiarities. Akin to the optical case, the gradient force , Eq. (22), plays the central role in acoustic tweezers. However, in the majority of optical systems the electric-dipole part of this force is relevant, while the magnetic-dipole part can be neglected. In contrast, in acoustic systems both the monopole and dipole parts of the gradient force are usually important. Moreover, these parts often have opposite directions, so the competition between the magnitudes of the monopole and dipole gradient forces essentially determines the trapping properties of the acoustic field.
Consider the simplest 1D standing sound wave:
@@
Evidently, the nodes of the pressure energy density correspond to the anti-nodes of the velocity energy density and vice versa. Therefore, when the monopole and dipole polarizabilities (real-valued, for simplicity) of the particle, and , have the same sign, the monopole and dipole parts of the gradient force have opposite directions (see also Section V.5).
@@
Akin to the first optical trapping experiments [36], the first acoustic tweezers were based on the standing-wave field formed by two counter-propagating beams [583]. Standing waves remain one of the main configurations used in acoustic and acoustofluidic manipulations [197, 26, 261, 550, 154, 185, 584, 340, 334, 412, 585, 375, 184, 62, 505], Fig. 17(f) (see also Section V.5). Curiously, the first historical experiment demonstrating mechanical action of acoustic waves was also implemented in a standing sound wave: in 1866 Kundt observed accumulation of powder particles near the standing-wave nodes [312, 472].
@@
More recently, single-beam acoustic tweezers came into play [325, 49], and these are often based on the use of vortex beams [49, 50, 48, 63, 61, 611, 369, 367, 368, 62]. Considering paraxial acoustic vortex beams (see Section V.2), the pressure and velocity energy densities have similar distribution, , with a nodal line along the beam’s axis, Fig. 17(c). In this case, a ‘negative’ trapping gradient force can be achieved for relatively hard particles in air. Indeed, typical particles are heavier but less compressible than air: , . According to Appendix B, this means that and . Thus, the dipole and monopole parts of the radial gradient force have opposite signs; the trapping in the vortex center is achieved when the monopole contribution prevails. Naturally, vortex-beam trapping often involves ‘orbital’ torques induced by the vortex.
@@
Akin to optics, holographic acoustic tweezers allowing arbitrary 2D and 3D arrangements and manipulation of multiple particles have been produced [373, 369, 368, 263], see Figs. 1(c), 3(b) and 17(g-j). Remarkably, holographic acoustic manipulations via phase-gradient, i.e., radiation-pressure force have also been demonstrated [373], Fig. 17(i-j).
@@
Finally, we note that acoustic manipulations are often accompanied by the phenomenon of acoustic streaming, i.e., steady flows induced by the dissipation of sound waves in the fluid [339]. Such streams themselves can be used as effective tools for acoustofluidic manipulation and acoustic-streaming tweezers [7, 363, 256, 412, 375, 271, 184, 62].

### v3

@@ V.1.2 Acoustics
These equations are similar to electromagnetic expressions (51) with and , i.e., without polarization degrees of freedom. Consequently, they result in analogous basic forces (24) and torque (25) or (28) acting on an isotropic Rayleigh particle: (i) the -directed gradient force, (ii) the -directed radiation-pressure force, and (iii) the -directed torque due to the transverse spin . Such forces and torque have been analyzed in [602, 15, 214, 506].
The recoil force (26) is produced by the complex kinetic momentum density . Unlike the optical case (52), this momentum has only and components, and, hence, the recoil force produces only small corrections to the dominant gradient and radiation-pressure forces. In acoustics, obtaining a recoil force with a nontrivial transverse direction requires either anisotropic particles [567] or Willis coupling [532].
The evanescent-wave example is valuable not only per se but also as a generic model for fields where the intensity and phase gradients are orthogonal. For instance, the interference of two plane waves with equal amplitudes and frequencies but different propagation directions produces a field with its mean wavevector (phase gradient) directed along, say, the -axis, and with periodic intensity variations along the orthogonal -axis. Such fields produce analogous local forces and torques, both in optics [72, 146, 4, 579] and acoustics [540, 108].
@@ V.2 Vortex beams
Let us consider another important class of structured wave fields: paraxial vortex beams. Such beams carrying orbital angular momentum (OAM) were introduced in optics in the early 1990s [69, 17]. Since then, they have found widespread applications in optics [16, 606, 71, 23, 422, 223, 665, 537], acoustics [273, 349, 263, 143], and quantum matter waves [98, 613, 99, 382, 338, 160, 387]. Independently of their physical nature, paraxial vortex beams are characterized by the scalar vortex factor , where are cylindrical coordinates, is the radial function which behaves as for , and is the integer topological charge (azimuthal quantum number) of the vortex.
@@ V.2.1 Optics
where is the same complex polarization parameter introduced earlier for evanescent waves, and we neglected the small longitudinal -components of the fields. Substituting the fields (55) into Eqs. (2)–(8) yields the energy, momentum, and spin density distributions:
@@
Here, the azimuthal component of the canonical momentum, , generates the OAM density , which looks similar to the spin but with the vortex charge replacing the helicity parameter .
@@
Substituting Eqs. (56) into the main force and torque Eqs. (24), (25), and (28), we find that an isotropic absorbing Rayleigh particle in a paraxial vortex beam experiences: (i) a radial gradient force, (ii) longitudinal and azimuthal radiation-pressure forces, and (iii) a longitudinal torque due to the spin. Consequently, if the particle’s -position is fixed (e.g., by a substrate), it can: (i) be trapped at a radial maximum or minimum of the field intensity by the gradient force (depending on the sign of the real part of its dipole polarizability), (ii) orbit azimuthally along a circle due to the azimuthal radiation-pressure force controlled by the vortex charge , and (iii) spin around the local -axis passing through the particle due to the torque controlled by the polarization helicity , see Fig. 16(a,b). The azimuthal orbiting and local spinning clearly demonstrate the qualitative difference between the orbital (extrinsic) and spin (intrinsic) angular momenta. Such dynamics of small particles in optical vortex beams was observed in [228, 229, 446, 236, 617, 237, 170]. Note that the -controlled spinning observed in these experiments originated from the particle’s anisotropy and the corresponding anisotropy-induced torque (31) rather than the absorption torque (28). Employing anisotropy instead of absorption is typically more efficient and avoids undesirable heating.
@@
Particles experience ‘negative’ gradient forces and are attracted to low-intensity regions (‘dark-field’ trapping), when the real part of their polarizability is negative. This occurs when the refractive index of the dielectric particle is lower than the index of the surrounding medium [229, 236] or at over-resonant frequencies near polarizability resonances (Fig. 10) [191, 673, 431]. Such negative gradient forces enable trapping of negative-polarizability particles at the zero-field centers of vortex beams [523, 560, 447, 255].
@@
When a particle occupies the central region of the vortex beam (e.g., when it is trapped at the center or its size exceeds the characteristic vortex radius), both the azimuthal radiation-pressure force and the spin-induced torque produce similar rotational effects with respect to the beam/particle center. In this regime, the ‘orbital’ and ‘spin’ torques act in similar ways [272, 559], and the OAM-controlled ‘orbital torque’ can be used for local rotational manipulation of particles [660, 35].
@@
Furthermore, beyond the paraxial-field approximation (55) and the basic gradient and radiation-pressure forces, an additional -dependent orbital rotation of particles can appear in vortex beams. This occurs via two mechanisms. First, in nonparaxial vortex beams, there is a -dependent contribution to the azimuthal canonical momentum (i.e., OAM ) [5, 686, 88, 109, 105]. Second, a similar effect can arise due to the -dependent azimuthal component of the complex Poynting (kinetic) momentum density and the associated recoil force (26) originating from its real [24, 25, 581] and imaginary [691, 654] parts.
@@ V.2.2 Acoustics
In acoustics, paraxial vortex beams can be described by the scalar pressure field and the corresponding velocity field . This longitudinal sound field produces the energy and canonical momenta densities and similar to electromagnetic Eqs. (56) but without spin density: . (The transverse spin, present in both optics and acoustics, arises only in nonparaxial fields that have both transverse and longitudinal components [105, 11, 106].) As in optics, the azimuthal component of the canonical momentum is responsible for the OAM carried by the beam.
Accordingly, paraxial acoustic vortex beams exert only a radial gradient force and a radiation pressure force with longitudinal and azimuthal components. A number of experiments have demonstrated acoustic ‘orbital torques’ on Mie-sized or larger particles at the center of acoustic vortex beams, driven by the azimuthal radiation-pressure force [618, 565, 179, 26, 53]. The dynamical effects of acoustic vortices have found numerous applications, including trapping, levitation, and holographic tweezers [401, 52, 282, 399, 400, 64], rotation of particles up to 6500 revolutions per second [676], manipulation of cells [66, 682], microbubbles [51] and objects in living bodies [241] (see [263] for a review).
Figure 16(c) shows experimental observations [281] of the local orbital motion of microparticles in acoustic vortex fields. This microparticle velocity field can be associated with the local canonical momentum density or Stokes drift distribution in acoustic vortices [see Section IV.10 and Fig. 14(b)].
@@ V.3.1 Optics
Optical tweezers have been the most prolific implementation of optical forces. They use one or more tightly focused laser beams to trap and manipulate particles. Optical trapping was first proposed and demonstrated by Ashkin and co-workers [45, 41, 44, 46]. The explanation of the effect depends on the size of the trapped object, Eq. (III.1). For particles larger than several wavelengths, a ray-optics picture is appropriate. Namely, large dielectric particles act as lenses, re-directing the momentum of the incident rays, with the corresponding recoil pushing them towards the focus [40, 41]. For particles much smaller than the wavelength, the dipolar regime dominates, and the intensity-gradient force in Eq. (24) pushes the particle toward the high-intensity focus. For particles of intermediate size, comparable to the wavelength, neither model is accurate, and one has to consider higher-order multipoles, Mie scattering calculations, or numerical calculations of forces based on the general Eqs. (14).
@@
Importantly, optical trapping can be applied to individual atoms [158, 466, 478], Bose-Einstein condensates [265, 155], biological objects [44, 46], and various nanostructures such as quantum dots, nanotubes, and graphene flakes [394, 456, 584, 299]. Optical tweezers can also be used in nanofabrication, direct microfluidic control, and assembly [230, 591, 456]. Beyond manipulation, they enable precise measurement of tiny particle displacements, since light scattering is highly sensitive to the local intensity profile within the focus.
@@
Let us consider the main field configurations used in optical tweezers for small particles. In the simplest case, a single focused beam is employed. Assuming electric-dipole particles, the trapping gradient force (for ) competes with destabilizing radiation-pressure and recoil forces and , which push the particle longitudinally out of the trap, Fig. 17(a). According to Sections III.3 and III.4, the gradient force depends on the tightness of focusing and on , while the pressure force is proportional to . For lossless Rayleigh particles with , we have , Eq. (III.3), so the gradient force typically dominates over the radiation-pressure one. However, both pressure and recoil forces grow with the particle size and refractive index faster than the gradient force, making it challenging to trap larger dielectric particles or high-index particles. This limitation can be overcome using very tight focusing and/or anti-reflection coatings [113, 298]. In addition, for small particles, the gradient force competes with Brownian motion, which depends strongly on the surrounding medium and temperature. This can be problematic for metallic (lossy) nanoparticles that heat up under illumination [531]. At the same time, polarizability resonances in metallic particles can be advantageously exploited for trapping and control [190, 459].
@@
The next field configuration involves two counter-propagating focused beams with equal amplitudes, polarizations, and other parameters. This arrangement produces a standing wave with alternating maxima and minima of the electric energy density, separated by half a wavelength along the longitudinal axis, Fig. 17(b). Importantly, in such a non-propagating field the momentum and energy flux densities vanish, and so do the pressure and recoil forces. Counter-propagating beams were first used in Ashkin’s pioneering work [39, 41] to trap large particles with sizes of several wavelengths. For such particles, the fine subwavelength standing-wave features are unimportant, and the wavelength-averaged intensity gradients provide stable trapping in the focal region. For subwavelength Rayleigh particles, the standing-wave structure becomes important, and such particles can be trapped at local field-intensity maxima (for ) [674] or minima () [673], as shown in Fig. 17(b). Standing-wave trapping is particularly relevant in the context of cavity optomechanics [48].
@@
Optical tweezers can also employ structured (essentially non-Gaussian) beams. In particular, both high-intensity (for ) and low-intensity (for ) regions have been exploited for trapping in vortex beams, see Section V.2, Fig. 16, and Fig. 17(c). Furthermore, “optical bottle beams”, which generate a 3D dark spot, have been used for ‘dark-field’ trapping [32, 293, 652, 57, 550]. Note that photophoretic forces can play an important role in such ‘dark-field’ trapping with vortex and bottle beams [549, 185, 551, 550, 568]. Diffraction-free Bessel beams [201] have been used to trap and manipulate particles over long distances in both single-beam [31] and counter-propagating two-beam [138, 314, 659] configurations. Likewise, Airy beams that exhibit ‘transverse acceleration’ upon propagation [82, 564] have also been applied in manipulation experiments [68, 67].
@@
A breakthrough in optical tweezers came with the introduction of multi-functional traps, often called holographic tweezers [261]. These systems generate arrays of multiple beams with independently controlled foci, enabling simultaneous multi-point manipulation of many trapped particles [199, 521, 421, 522, 200, 490, 367, 171, 56, 568], see Figs. 1(b), 3(a), and 17(e).
@@
Importantly, optical trapping can be combined with various torques, enabling controlled rotations of anisotropic or absorptive trapped particles. This has driven rapid progress in the study of optical torques and their applications [226, 272, 559, 451, 127]. Remarkably, the achievable angular velocities of trapped particles span an enormous range: from quasi-static 3D alignments and slow (Hz-scale) rotations [459, 604, 647] to MHz–GHz rotations, where material strength limits and quantum effects become relevant [30, 330, 287, 9, 491, 254].
@@
Optical trapping and tweezers have found wide-ranging applications across many fields. First, they have become an invaluable tool for manipulating and probing biomolecules and living cells [44, 46, 111, 583, 1, 38, 571, 582, 188, 677, 127, 650], Fig. 17(d). Second, optical trapping has played a key role in optical cooling of atoms, molecules, nanoparticles, and larger macroscopic objects [158, 466, 412, 141, 361, 246, 166, 587, 360, 177, 524, 509], enabling advances from macroscopic quantum interference [33, 505] to gravitational-wave detection [34]. Third, optical forces near the trap centers create effective harmonic-oscillator potentials, providing a platform for both classical and quantum optomechanical studies [48, 413, 142, 254, 297, 248, 588]. Moreover, 1D, 2D, and 3D arrays of optically trapped atoms, molecules, or particles can serve for quantum simulators and computers [178, 210, 55, 78, 630, 526, 203, 56], atomic clocks [669], and volumetric displays [568], see Figs. 3(a) and 17(e). Finally, optical trapping and tweezers play the key role in various force-microscopy techniques [433, 575].
@@ V.3.2 Acoustics
Acoustic trapping and tweezers share natural similarities with their optical counterparts, but also exhibit important peculiarities. As in optics, the gradient force , Eq. (24), plays the central role in acoustic tweezers. However, in most optical systems the electric-dipole part of this force dominates, while the magnetic-dipole part can typically be neglected. In contrast, in acoustic systems, both the monopole and dipole parts of the gradient force are usually significant. Moreover, these parts often act in opposite directions, so the competition between the magnitudes of the monopole and dipole gradient forces largely determines the trapping properties of the acoustic field.
Consider the simplest case of a 1D standing sound wave:
@@
Here, the nodes of the pressure energy density coincide with the anti-nodes of the velocity energy density , and vice versa. Therefore, when the monopole and dipole polarizabilities of the particle, and (real-valued, for simplicity), have the same sign, the monopole and dipole parts of the gradient force act in opposite directions (see also Section V.5).
@@
Similarly to the first optical trapping experiments [39], the earliest acoustic tweezers were based on a standing-wave field formed by two counter-propagating beams [643]. Standing waves remain one of the primary configurations for acoustic and acoustofluidic manipulation [207, 28, 276, 609, 163, 194, 645, 369, 359, 450, 646, 410, 193, 65, 558], Fig. 17(f) (see also Section V.5). Interestingly, the first historical demonstration of the mechanical action of acoustic waves was also implemented in a standing sound wave: in 1866 Kundt observed the accumulation of powder particles near the standing-wave nodes [333, 520].
@@
More recently, single-beam acoustic tweezers have been realized [346, 52], often employing vortex beams [52, 53, 51, 66, 64, 676, 401, 399, 400, 65]. For paraxial acoustic vortex beams (see Section V.2), the pressure and velocity energy densities have similar distributions, , with a nodal line along the beam’s axis, Fig. 17(c). In this case, a ‘negative’ trapping gradient force can be achieved for relatively hard particles in air. Indeed, typical particles are heavier but less compressible than air: , . According to Appendix B, this means that and . Thus, the dipole and monopole parts of the radial gradient force have opposite signs; trapping in the vortex center occurs when the monopole contribution prevails. Naturally, vortex-beam trapping also involves ‘orbital’ torques induced by the vortex phase structure.
@@
As in optics, holographic acoustic tweezers have been developed, enabling arbitrary 2D and 3D arrangements and simultaneous manipulation of multiple particles [408, 401, 400, 278], see Figs. 1(c), 3(b) and 17(g-j). Notably, holographic acoustic manipulation via phase-gradient, i.e., using the radiation-pressure force , has also been achieved [408], Fig. 17(i-j).
@@
Finally, acoustic manipulation is often accompanied by acoustic streaming: steady fluid flows induced by the dissipation of sound waves [368]. Such flows themselves can serve as an efficient tool for acoustofluidic manipulation and acoustic-streaming tweezers [8, 395, 270, 450, 410, 286, 193, 65].
Numerous applications of acoustic trapping and tweezers partially parallel those of optical tweezers, including the manipulation of particles, microfluids, biological cells, and living organisms [450, 188, 211, 370, 410], noninvasive manipulation in living bodies [241, 664, 308], holographic tweezers and volumetric displays [401, 400, 278], contactless manipulation and probing of extraterrestrial materials [220], and acoustic force spectroscopy [563].